Systems Staff Meeting Notes 4/16/01

Systems Staff Meeting Notes
April 16, 2001

Present: Cynthia Bowker, Will Chaney, Bill Covey, Debra Harris, Suzy Shaw, Hilton Soto, Geoffrey Stewart, Ying Tang
Announcements
  1. Will Chaney vacationed in Daytona over the weekend.
  2. W2K Training was intense last week. The general consensus is that a central tree will not work in the immediate future. We’ll have to go with individual trees. IRC report was also a hot topic.
Ongoing Systems Updates
Equipment Received/Installed/Repaired
  1. Two Dell Precision 420 workstations for the Facilities Security system have been ordered. They are converting over to the Axis security system. Rob Roberts wants two BellStor raid units on the workstations. They may come to us, but Semco is supposed to install them. We’ll have to supply the IPs and should contact Digital regarding the network drops. Should also check on power.
  2. Laptops for MSL classroom still on hold.
  3. Security desk for room 100 East has been received. We’ll tell Gary in HS&S to have it assembled.
Liaisons
  1. Geoffrey notes that the new Dells are causing a lot of confusion for people. The BIOS doesn’t look like it used to – there are more Dells with different configurations in this batch. Might need to refresh the liaisons on basic procedures – perhaps it’s time for another orientation session?
Networking
  1. The Squid server is active again. The function was moved back to the box in Systems, so we can retrieve the one in West.
  2. Password changes have been made on all switches and servers following Bobbie’s departure from Systems.
  3. We’ll need to have the vault combinations changed as well – Debra will contact Facilities and get the process started.
  4. West closet AC – have not heard anything. Still waiting.
  5. Electric for the new Systems server room was started last week. It is scheduled to be complete by April 28. The rack guys are supposed to show on May 7. The Dell server guy should be here on May 8. We may need a server license – possibly two of them. Will will look into it.
  6. There was a lot of Internet trouble last week with the Internet POPs. The 3rd POP is scheduled to go in this summer.
Software Projects Status
  1. OPS Tracker – tested on Friday, some improvements are necessary – should be ready in a few days.
  2. IRDB – Winston is working on it.
  3. Personnel – again, Winston’s puppy.
  4. Elected Officials database – Cynthia met with Tom Minton & Trish Ruwell last week. Project should be finished this week.
  5. Ringling – Ying reviewed the project with Erich & Stephanie. They really liked the Web search feature. Ying needs to make a few final changes and add data to the database – then it will go to testing and finally production.
  6. Ying is back working on the Tracking database using DELPHI.
  7. ILLIAD – still no word.
DP Coordinator
  1. Students start finals next week.
  2. Debra is searching for new students for the summer. None of our current batch will be here, although two will return in the fall.
  3. Tape migration – Winston generated a list of empty rack/slots to begin the move. Debra will have the students start the process this week. Will try to have the first row on the west side of the room empty by the time the rack and server installation is scheduled to occur (May 7).
  4. Keep up with the calendar – Bill is out Wed-Fri this week.


Back to the Systems Homepage

Last updated April 30, 2001
by Debra Harris
debharr@mail.uflib.ufl.edu